# compat=1.1.0

### v1

@@ Abstract
Introduction– Landau’s Fermi liquid theory has long served as the foundational framework for understanding conventional metals, wherein low-energy excitations are described by long-lived quasiparticles with renormalized effective masses. In this picture, electrical resistivity saturates at both low and high temperatures—due to phonon freezing and the Mott–Ioffe–Regel (MIR) bound, respectively [1, 2]. However, in a growing class of materials known as strange metals, the resistivity remains linear in temperature over a remarkably wide range without saturation, even beyond the MIR limit, indicating the breakdown of the quasiparticle description and the emergence of collective, nonlocal charge-carrying degrees of freedom [3, 4, 5, 6, 7, 8]. This non-Fermi liquid behavior in strange metals is typically characterized by linear-in-temperature ($T$-linear) resistivity and linear-in-magnetic-field ($B$-linear) magnetoresistivity, alongside scale-invariant dissipation governed by the Planckian time $\tau_{P}=\hbar/(k_{B}T)$ [9, 11, 12, 10, 13, 14]. Interestingly, the Hall angle in the cuprates exhibits a distinct scaling $\theta_{H}\sim 1/T^{2}$, resembling Fermi-liquid-like behavior and therefore standing in sharp contrast to the ubiquitous $T$-linear resistivity of strange metals [9, 10]. Experimental studies — notably the early work by Chien et al. — observed such Hall-angle scaling and its sensitivity to impurity scattering, underscoring the inadequacy of single-relaxation-time transport models and pointing instead to multiple, possibly channel-dependent, scattering mechanisms underlying strange-metal phenomenology [15].
@@
From this viewpoint, the standard Fermi–Dirac distribution—and the corresponding Fermi liquid behavior—is analogous to the Hawking saddle: a disconnected, local description valid in regimes where quasiparticles dominate. Allowing the replica index $q$ to deviate slightly from unity, $q=1+\delta(T)$ with $\delta(T)=T/T_{0}\ll 1$ temperature dependent, yields a Tsallis-deformed fermionic distribution that encodes non-Fermi liquid behavior, reminiscent of the contribution of replica wormholes in semiclassical gravity. Such a deformation can be captured by a Schwarzian effective action [27, 28], which governs the soft mode dynamics of nearly-$\mathrm{AdS}_{2}$ geometries. This correspondence suggests a dual statistical picture: the strange metal phase, with its scale-invariant transport and Planckian dissipation, may be understood as the condensed matter analogue of gravitational systems dominated by wormhole saddles. In the following, we develop this idea by deriving the quantum Tsallis distribution, relating it to both Rényi and von Neumann entropies, and showing how its low-energy limit maps to the Schwarzian theory that appears in the strongly interacting quantum matter.
@@
[…]
@@
$\mathcal{L}=\frac{1}{1-q}\ln\bigg({\rm Tr}\rho^{q}\bigg{missing})-\alpha({\rm Tr}\rho-1)-\beta\bigg{(}{\rm Tr}(\rho H)-\langle E\rangle\bigg{)}.$ (4)
@@
$\delta\mathcal{L}=\frac{q}{1-q}\frac{{\rm Tr}(\rho^{q-1}\delta\rho)}{{\rm Tr}(\rho^{q})}-\alpha{\rm Tr}\delta\rho-\beta{\rm Tr}(H\delta\rho).$ (5)
From maximal entropy $\delta\mathcal{L}=0$ and after some manipulation, we find
@@
which is finite whenever the trace exists; for unbounded $H$, one either has a high-energy cutoff for $q<1$ (since $1-(1-q)\beta\varepsilon\geq 0$ implies $\varepsilon\leq 1/[(1-q)\beta]$) or a power-law tail for $q>1$ with convergence controlled by the growth of the spectral density. In the Gibbs limit $q\to 1$, the well-known identity $\lim_{q\to 1}\big{[}1-(1-q)\beta H\big{]}^{1/(q-1)}=e^{-\beta H}$ applies eigenvalue-wise and gives $\rho\;\to\;\frac{e^{-\beta H}}{{\rm Tr}(e^{-\beta H})},\qquad Z_{q}\;\to\;Z={\rm Tr}(e^{-\beta H}),$ so the Tsallis ensemble continuously deforms the canonical ensemble and recovers it at $q=1$.
For quantum statistics with indistinguishable particles, the same $q$-deformation propagates to the grand-canonical occupation numbers once one replaces the ordinary exponential by the $q$-exponential while enforcing the exchange symmetry. The generalized quantum occupations can be written compactly as
$n_{q}(\varepsilon)\;=\;\frac{1}{\big{[}\,1-(1-q)\beta(\varepsilon-\mu)\,\big{]}_{+}^{\frac{1}{1-q}}+a},$
where $\beta=1/T$ and $\mu$ is the chemical potential; the parameter $a=+1$ gives the Tsallis–Fermi–Dirac case (Pauli exclusion), a=1a=-1 the Tsallis–Bose–Einstein case (Bose enhancement), and $a=0$ the Tsallis–Boltzmann classical limit. This expression reduces to the standard Fermi–Dirac/Bose–Einstein/Boltzmann forms as $q\to 1$. If one prefers the alternative but equivalent representation used in some conventions, the denominator can be written as $\big{[}1+(q-1)\beta(\varepsilon-\mu)\big{]}^{1/(q-1)}+a$, noting that $1+(1-q)x]^{1/(q-1)}=\big{(}[1+(q-1)(-x)]^{1/(q-1)}\big{)}^{-1}$; either choice is consistent with the operator formula for $\rho$ above, provided the same convention for the $q$-exponential is used throughout and the spectral positive-part enforces positivity and, when $q<1$, the finite-energy support of the ensemble.
@@
Thus, $\phi(x)$ can be interpreted as the variance of the Fermi–Dirac occupation number, modulated by a quadratic prefactor $x^{2}$. Since $n_{\mathrm{FD}}(1-n_{\mathrm{FD}})$ is maximized at $x=0$ and decays away from the Fermi surface, this thermal weight confines the correction to states near the Fermi level. In the frequency–time correspondence (taking the Euclidean Matsubara convention), we have $x=\beta\omega$ and $i\partial_{\tau}\leftrightarrow\omega$. Hence,
@@
$S_{q}[f]=-C\int\{f,\tau\}\,+\,(q-1)\,\lambda\int\Phi[f](\tau)\,+\,\mathcal{O}((q-1)^{2}),$
where $\{f,\tau\}$ is the Schwarzian derivative and $\lambda$ is a coupling determined by the microscopic model. For small deviations from the identity reparametrization, $f(\tau)=\tau+\varepsilon(\tau)$, the quadratic effective action for the soft-mode fluctuations takes the form
$S_{q}^{(2)}[\varepsilon]=\frac{\bar{C}_{\rm eff}}{2}\int_{0}^{\beta}d\tau\,(\varepsilon^{\prime\prime}(\tau))^{2},$ (9)
@@
$\bar{C}_{\rm eff}=\bar{C}+\delta C\,(q-1)+\mathcal{O}((q-1)^{2}).$
In the limit $q\to 1$, one recovers the standard Schwarzian coupling $\bar{C}$ and the familiar zero-mode solutions
@@
$Z_{q}=\int\frac{\mathcal{D}f}{SL(2,\mathbb{R})}\exp\left\{-\bar{C}\int_{0}^{\beta}d\tau\left[\{f(\tau),\tau\}-(q-1)\left(\frac{f^{\prime\prime}(\tau)}{f^{\prime}(\tau)}\right)^{2}\right]\right\},$ (10)
where $\bar{C}$ is the Schwarzian coupling constant. The $q$-deformation of Eq. (10) contributes an additional quadratic term to the effective action, which can be interpreted as a renormalization of the Schwarzian stiffness and modifies the dynamics of the soft reparametrization mode. The corresponding effective Hamiltonian reads
$H=-\frac{1}{2\bar{C}}\frac{d^{2}}{du^{2}}+\frac{1}{8\bar{C}\sinh^{2}(u)},$ (11)
@@
The spectral density is determined by the Plancherel measure of $SL(2,\mathbb{R})$ representations  [27, 28, 31, 32, 33]
$\rho(E)=\frac{\bar{C}}{2\pi^{2}}\sinh\!\left(2\pi\sqrt{2\bar{C}E}\right),$ (14)
@@
$\rho(E)\approx\frac{\bar{C}^{3/2}}{\pi}\sqrt{2E},~E\ll 1;~~~\rho(E)\sim\frac{\bar{C}}{4\pi^{2}}\exp\!\left(2\pi\sqrt{2\bar{C}E}\right),~E\gg 1.$ (15)
@@
$S^{(2)}_{\rm Sch}=2\pi\bar{C}T\sum_{n\geq 2}n(n^{2}-1)\,|\varepsilon_{n}|^{2},\qquad\varepsilon_{-n}=\varepsilon_{n}^{*}.$ (16)
@@
$\Delta S^{(2)}_{q}=2\pi\bar{C}T\,\gamma_{q}\sum_{n\geq 2}n^{2}\,|\varepsilon_{n}|^{2},\qquad\gamma_{q}=\alpha\,(q-1),$ (17)
@@
$S^{(2)}_{\rm tot}=2\pi\bar{C}T\sum_{n\geq 2}\Big{[}n(n^{2}-1)+\gamma_{q}n^{2}\Big{]}\,|\varepsilon_{n}|^{2},$ (18)
where the first term $n(n^{2}-1)$ is the standard Schwarzian spectrum after fixing the $SL(2,\mathbb{R})$ zero modes, while the second term $\gamma_{q}n^{2}$ represents the leading $q$-deformation effect. Physically, the additive $\gamma_{q}n^{2}$ term modifies the dispersion of the soft reparametrization modes. While the Schwarzian spectrum $n(n^{2}-1)$ grows cubically with $n$, the $q$-deformation introduces a quadratic contribution that enhances the stiffness of low-$n$ modes, most prominently the $n=2$ component. As a result, the $q$-deformation shifts the balance between universal Schwarzian fluctuations and non-Gaussian corrections inherited from the microscopic dynamics. The Matsubara propagator for the reparametrization (soft) modes is defined as the frequency-space two-point function
@@
$D(i\omega_{n})=\langle|\varepsilon_{n}|^{2}\rangle=\frac{1}{2\pi\bar{C}T}\,\frac{1}{\,A_{n}\,}=\frac{1}{2\pi\bar{C}T}\;\frac{1}{n(n^{2}-1)+\gamma_{q}n^{2}}.$ (20)
@@
with a non-universal prefactor $\kappa_{b}=\mathcal{O}(1)$. The deformation introduces an effective scale
$T_{0}\equiv\frac{\alpha}{2\pi}(q-1)\Lambda\;\;\propto\;\;\gamma_{q}T,$ (22)
where $\Lambda\sim 2\pi T$ is the thermal UV cutoff. The emergent Tsallis scale $T_{0}$ is quantitatively anchored by the Raman low-frequency slope, whose ratios across doping levels yield consistent values around $100\text{K}$, as shown in the Appendix. Equation (21) shows that the $q$-deformation regularizes the infrared behavior of the soft modes by shifting the effective bath temperature to $T+T_{0}$. In the undeformed limit $q\to 1$, one recovers the canonical Schwarzian bath with Ohmic spectral density $A_{\varepsilon}(\omega)\propto|\omega|/T$ [39].
@@
$K(\omega)=\frac{\kappa}{\bar{C}}\,\frac{|\omega|}{T_{0}+T},\qquad\kappa=\frac{g^{2}\kappa_{b}}{2\pi},$ (29)
which makes explicit that the effective interaction strength scales inversely with the renormalized Schwarzian stiffness $\bar{C}$.
After integration, we obtain the compact form
@@
$\sigma_{xx}\simeq e^{2}\int d\epsilon\,(-\partial_{\epsilon}f)\,N(\epsilon)\,v_{x}^{2}(\epsilon)\,\tau_{\rm tr}(\epsilon),\qquad\tau_{\rm tr}(\epsilon)\sim\Gamma^{-1}(\epsilon),$ (31)
@@
Hall angle from the memory matrix- The memory-matrix formalism provides a controlled framework for transport in systems with slow or nearly conserved operators, where standard Boltzmann theory fails [48, 49, 50]. The central idea is that dc conductivities are determined by the relaxation of such slow operators through their couplings to a dissipative bath. For strange metals, this bath is supplied by Schwarzian soft modes, whose Ohmic spectrum ensures smooth low-frequency behavior and introduces an additional channel for energy dissipation. Concretely, for a basis of slow operators $\{A_{\alpha}\}$ the conductivity tensor reads
$\sigma_{ij}=\sum_{\alpha\beta}\chi_{J_{i}A_{\alpha}}\,\big{[}M(0)+N\big{]}^{-1}_{\alpha\beta}\,\chi_{A_{\beta}J_{j}},$ (33)
where $M$ encode the dissipative couplings to the bath, $\chi_{\alpha\beta}=\langle A_{\alpha};A_{\beta}\rangle$ the static susceptibilities. This expression applies equally to Fermi liquids and non–Fermi liquids. In the Fermi-liquid case, $M$ reduces to the standard quasiparticle collision integral, recovering the textbook Kubo–bubble results. $N$ is the antisymmetric block induced by the Lorentz force, with $N_{P_{x}P_{y}}=-N_{P_{y}P_{x}}=\omega_{c}\chi_{PP}$ and $\omega_{c}=eB/m^{\ast}$. For completeness, we rederive the dc conductivity and the quadratic magnetoresistance of a conventional Fermi liquid in Appendices A and B. The key quantities are the relaxation rates
$\Gamma_{Q}\equiv\frac{M_{QQ}(0)}{\chi_{QQ}},\qquad\Gamma_{P}\equiv\frac{M_{PP}(0)}{\chi_{PP}},$
which characterize dissipation in the energy-like channel $Q$ and in the momentum channel $P$, respectively. When $Q$ is chosen as the current-like soft mode, Eq. (33) reduces to
$\sigma_{xx}=\frac{\chi_{JJ}}{\Gamma_{Q}},\qquad\rho_{xx}=\frac{\Gamma_{Q}}{\chi_{JJ}}.$ (34)
@@
$M_{QQ}(0)\propto T\qquad\Rightarrow\qquad\rho_{xx}\propto T.$ (35)
Physically, the soft reparametrization modes relax energy efficiently without relaxing total momentum, giving rise to a Planckian-like linear resistivity channel. For the momentum block,
@@
In a continuum model one has $\dot{P}_{x}=0$, so momentum relaxation arises solely from translation-breaking terms such as static disorder $V(\mathbf{r})$ and lattice umklapp. For a weak periodic potential $V(\mathbf{r})=\sum_{\mathbf{G}}U_{\mathbf{G}}e^{i\mathbf{G}\cdot\mathbf{r}}$, the coefficients $U_{\mathbf{G}}$ are Fourier components of the lattice potential at reciprocal vectors $\mathbf{G}$. The operator $\rho_{\mathbf{G}}=\sum_{\mathbf{k},\sigma}c^{\dagger}_{\mathbf{k}+\mathbf{G},\sigma}c_{\mathbf{k},\sigma}$ transfers electronic density by momentum $\mathbf{G}$, while $G_{x}$ denotes the $x$-component of $\mathbf{G}$, quantifying the amount of momentum lost along the transport direction. The leading memory-matrix element controlling momentum relaxation then reads
[…]
where $\chi^{R}$ is the retarded density–density correlator of the interacting system. In clean correlated metals the dominant temperature dependence arises from electron–electron umklapp scattering at finite $\mathbf{G}$. Phase-space integrals over two quasiparticles near the Fermi surface yield the familiar Fermi-liquid factor
$\int d\epsilon_{1}d\epsilon_{2}\,\epsilon_{1}\epsilon_{2}\,(-\partial_{\epsilon_{1}}f)(-\partial_{\epsilon_{2}}f)\;\propto\;T^{2},$ (38)
so that
$M_{PP}(0)\propto T^{2},\qquad\Gamma_{P}=\frac{M_{PP}}{\chi_{PP}}\propto T^{2}.$ (39)
Hence momentum relaxation is parametrically slower than energy relaxation, consistent with conventional large-angle umklapp scattering. Because $Q$ couples strongly to the current but weakly to the Lorentz-force channel, while $P$ couples directly to the Lorentz block $N$, the transport coefficients separate naturally:
$\rho_{xx}\propto\Gamma_{Q}\sim T,\qquad\cot\theta_{H}\simeq\frac{M_{PP}}{\omega_{c}\chi_{PP}}\propto T^{2}.$ (40)
@@
This dichotomy encapsulates the two-lifetime phenomenology of strange metals. Deviations from $\cot\theta_{H}\propto T^{2}$ in electron-doped cuprates [51] and ruthenates [52] reflect orbital selectivity and multi-band scattering, but the separation of channels remains a robust organizing principle [53].

### v2

# From Quantum Tsallis Entropy to Strange Metals

@@ Abstract
Introduction– Landau’s Fermi liquid theory has long served as the foundational framework for understanding conventional metals, wherein low-energy excitations are described by long-lived quasiparticles with renormalized effective masses. In this picture, electrical resistivity saturates at both low and high temperatures—due to phonon freezing and the Mott–Ioffe–Regel (MIR) bound, respectively [1, 2]. However, in a growing class of materials known as strange metals, the resistivity remains linear in temperature over a remarkably wide range without saturation, even beyond the MIR limit, indicating the breakdown of the quasiparticle description and the emergence of collective, nonlocal charge-carrying degrees of freedom [3, 4, 5, 6, 7, 8]. This non-Fermi liquid behavior in strange metals is typically characterized by linear-in-temperature ($T$-linear) resistivity and linear-in-magnetic-field ($B$-linear) magnetoresistivity, alongside scale-invariant dissipation governed by the Planckian time $\tau_{P}=\hbar/(k_{B}T)$ [9, 11, 12, 10, 13, 14]. Interestingly, the Hall angle in the cuprates exhibits a distinct scaling $\cot\theta_{H}\sim T^{2}$, resembling Fermi-liquid-like behavior and therefore standing in sharp contrast to the ubiquitous $T$-linear resistivity of strange metals [9, 10]. Experimental studies — notably the early work by Chien et al. — observed such Hall-angle scaling and its sensitivity to impurity scattering, underscoring the inadequacy of single-relaxation-time transport models and pointing instead to multiple, possibly channel-dependent, scattering mechanisms underlying strange-metal phenomenology [15].
@@
From this viewpoint, the standard Fermi–Dirac distribution, and the corresponding Fermi liquid behavior, is analogous to the Hawking saddle: a disconnected, local description valid in regimes where quasiparticles dominate. Allowing the replica index $q$ to deviate slightly from unity, $q=1+\delta(T)$ with $\delta(T)=T/T_{0}<1$ temperature dependent, yields a Tsallis-deformed fermionic distribution that encodes non-Fermi liquid behavior, reminiscent of the contribution of replica wormholes in semiclassical gravity. Such a deformation can be captured by a Schwarzian effective action [27, 28], which governs the soft mode dynamics of nearly-$\mathrm{AdS}_{2}$ geometries. This correspondence suggests a dual statistical picture: the strange metal phase, with its scale-invariant transport and Planckian dissipation, may be understood as the condensed matter analogue of gravitational systems dominated by wormhole saddles. In the following, we develop this idea by deriving the quantum Tsallis distribution, relating it to both Rényi and von Neumann entropies, and showing how its low-energy limit maps to the Schwarzian theory that appears in the strongly interacting quantum matter.
@@
$S_{q}^{\mathrm{Tsallis}}=\frac{1}{q-1}\left[1-e^{(1-q)S_{q}^{\mathrm{Rényi}}}\right].$ (3)
@@
$\mathcal{L}=\frac{1}{1-q}\ln\bigg({\rm Tr}\rho^{q}\bigg.)-\alpha({\rm Tr}\rho-1)-\beta\bigg({\rm Tr}(\rho H)-\langle E\rangle\bigg).$ (4)
@@
$\delta\mathcal{L}=\frac{q}{1-q}\frac{{\rm Tr}(\rho^{q-1}\delta\rho)}{{\rm Tr}(\rho^{q})}-\alpha{\rm Tr}\delta\rho-\beta{\rm Tr}(H\delta\rho),$ (5)
where $\alpha=\mu/T$ and $\beta=1/T$. From maximal entropy $\delta\mathcal{L}=0$ and after some manipulation, we find
@@
which is finite whenever the trace exists; for unbounded $H$, one either has a high-energy cutoff for $q<1$ (since $1-(1-q)\beta\varepsilon\geq 0$ implies $\varepsilon\leq 1/[(1-q)\beta]$) or a power-law tail for $q>1$ with convergence controlled by the growth of the spectral density. For quantum statistics with indistinguishable particles, the same $q$-deformation propagates to the grand-canonical occupation numbers once one replaces the ordinary exponential by the $q$-exponential while enforcing the exchange symmetry. The generalized quantum occupations can be written compactly as
$n_{q}(\varepsilon)\;=\;\frac{1}{\big[\,1-(1-q)\beta(\varepsilon-\mu)\,\big]_{+}^{\frac{1}{q-1}}+a},$
where $\beta=1/T$ and $\mu$ is the chemical potential; the parameter $a=+1$ gives the Tsallis–Fermi–Dirac case (Pauli exclusion), a=1a=-1 the Tsallis–Bose–Einstein case (Bose enhancement), and $a=0$ the Tsallis–Boltzmann classical limit. This expression reduces to the standard Fermi–Dirac/Bose–Einstein forms as $q\to 1$. Using the alternative convention commonly employed in the literature, the denominator may be written as [ $\big[\,1+(q-1)\beta(\varepsilon-\mu)\,\big]^{\frac{1}{q-1}}+a.$ The deformation parameter $q$ organizes the ensemble into two regimes related by the thermodynamic duality $q\leftrightarrow(2-q)$. For $q>1$, the distribution exhibits non-extensive algebraic tails and unbounded energy support, while $q<1$ yields a compact-support ensemble determined by the spectral cutoff. In strange metals we focus on $q>1$, consistent with the observed long-tailed responses.
@@
Thus, $\phi(x)$ can be interpreted as the variance of the Fermi–Dirac (FD) occupation number, modulated by a quadratic prefactor $x^{2}$. Since $n_{\mathrm{FD}}(1-n_{\mathrm{FD}})$ is maximized at $x=0$ and decays away from the Fermi surface, this thermal weight confines the correction to states near the Fermi level. In the frequency–time correspondence (taking the Euclidean Matsubara convention), we have $x=\beta\omega$ and $i\partial_{\tau}\leftrightarrow\omega$. Hence,
@@
$S_{q}[f]=S_{sch}+\Delta S_{q}=-C\int\{f,\tau\}\,+\,(q-1)\,\lambda\int\Phi[f](\tau)\,+\,\mathcal{O}((q-1)^{2}),$ (9)
where $\{f,\tau\}$ is the Schwarzian derivative and $\lambda$ is a coupling determined by the microscopic model. For small deviations from identity reparametrization, $f(\tau)=\tau+\varepsilon(\tau)$, the effective quadratic action for soft-mode fluctuations takes the form of
$S_{q}[\varepsilon]=\frac{\bar{C}_{\rm eff}}{2}\int_{0}^{\beta}d\tau\,(\varepsilon^{\prime\prime}(\tau))^{2},$ (10)
@@
$\bar{C}_{\rm eff}=C+\delta C\,(q-1)+\mathcal{O}((q-1)^{2}).$
In the limit $q\to 1$, one recovers the standard Schwarzian coupling $C$ and the familiar zero-mode solutions
@@
$Z_{q}=\int\frac{\mathcal{D}f}{SL(2,\mathbb{R})}\exp\left\{-{C}\int_{0}^{\beta}d\tau\left[\{f(\tau),\tau\}-(q-1)\left(\frac{f^{\prime\prime}(\tau)}{f^{\prime}(\tau)}\right)^{2}\right]\right\},$ (11)
where C{C} is the Schwarzian coupling constant. The $q$-deformation of Eq. (11) contributes an additional quadratic term to the effective action, which can be interpreted as a renormalization of the Schwarzian stiffness and modifies the dynamics of the soft reparametrization mode. The corresponding effective Hamiltonian reads
$H=-\frac{1}{2{C}}\frac{d^{2}}{du^{2}}+\frac{1}{8{C}\sinh^{2}(u)},$ (12)
@@
where $P_{\nu}^{\mu}(z)$ is the associated Legendre function of the first kind. The spectral density is determined by the Plancherel measure of $SL(2,\mathbb{R})$ representations [27, 28, 31, 32, 33]
$\rho(E)=\frac{{C}}{2\pi^{2}}\sinh\!\left(2\pi\sqrt{2{C}E}\right),$ (15)
@@
$\rho(E)\approx\frac{{C}^{3/2}}{\pi}\sqrt{2E},~E\ll 1;~~~\rho(E)\sim\frac{{C}}{4\pi^{2}}\exp\!\left(2\pi\sqrt{2{C}E}\right),~E\gg 1.$ (16)
@@
$S_{\rm Sch}=2\pi CT\sum_{n\geq 2}n(n^{2}-1)\,|\varepsilon_{n}|^{2},\qquad\varepsilon_{-n}=\varepsilon_{n}^{*}.$ (17)
@@
$\Delta S_{q}=2\pi CT\,\gamma_{q}\sum_{n\geq 2}n^{2}\,|\varepsilon_{n}|^{2},\qquad\gamma_{q}=\alpha\,(q-1),$ (18)
@@
$S_{\rm q}=2\pi CT\sum_{n\geq 2}\Big[n(n^{2}-1)+\gamma_{q}n^{2}\Big]\,|\varepsilon_{n}|^{2},$ (19)
where the first term $n(n^{2}-1)$ is the standard Schwarzian spectrum after fixing the $SL(2,\mathbb{R})$ zero modes, while the second term $\gamma_{q}n^{2}$ represents the leading $q$-deformation effect. Physically, the additive $\gamma_{q}n^{2}$ term modifies the dispersion of the soft reparametrization modes. While the Schwarzian spectrum $n(n^{2}-1)$ grows cubically with $n$, the $q$-deformation introduces a quadratic contribution that enhances the stiffness of low-$n$ modes, most prominently the $n=2$ component. As a result, the $q$-deformation shifts the balance between universal Schwarzian fluctuations and non-Gaussian corrections inherited from the microscopic dynamics. The Matsubara propagator for the reparametrization soft modes is defined as the frequency-space two-point function
@@
$D(i\omega_{n})=\langle|\varepsilon_{n}|^{2}\rangle=\frac{1}{2\pi CT}\,\frac{1}{\,A_{n}\,}=\frac{1}{2\pi CT}\;\frac{1}{n(n^{2}-1)+\gamma_{q}n^{2}}.$ (21)
@@
with a non-universal prefactor $\kappa_{b}=\mathcal{O}(1)$. The emergent Tsallis scale $T_{0}$ is quantitatively anchored by the Raman low-frequency slope, whose ratios across doping levels yield consistent values around $100\text{K}$, as shown in the Appendix C. Equation (22) shows that the $q$-deformation regularizes the infrared behavior of the soft modes by shifting the effective bath temperature to $T+T_{0}$. In the undeformed limit $q\to 1$, one recovers the canonical Schwarzian bath with Ohmic spectral density $A_{\varepsilon}(\omega)\propto|\omega|/T$ [39].
@@
$K(\omega)=\frac{\kappa}{C}\,\frac{|\omega|}{T_{0}+T},\qquad\kappa=\frac{g^{2}\kappa_{b}}{2\pi},$ (29)
which makes explicit that the effective interaction strength scales inversely with the renormalized Schwarzian stiffness $C$. The Fermi golden rule applies because the transition rate is universally determined by the product of the interaction matrix element and the available spectral phase space, regardless of whether the excitations form well-defined quasiparticles  [50]. After integration, we obtain the compact form
@@
$\sigma_{xx}\simeq e^{2}\int d\epsilon\,(-\partial_{\epsilon}n_{FD})\,N(\epsilon)\,v_{x}^{2}(\epsilon)\,\tau_{\rm tr}(\epsilon),\qquad\tau_{\rm tr}(\epsilon)\sim\Gamma^{-1}(\epsilon),$ (31)
@@
Hall angle from the memory matrix— The memory-matrix formalism, rooted in the Mori–Zwanzig projection-operator framework [48, 49], provides a controlled, non-perturbative scheme for transport in strongly interacting systems with nearly conserved slow operators, where quasiparticle-based Boltzmann theory fails [50]. The central idea is that dc transport is determined not by single-particle lifetimes but by the relaxation of conserved densities under weak perturbations.
In strange metals, the dominant incoherent relaxation channel arises from emergent Schwarzian soft modes—the Goldstone modes of the near-conformal time-reparametrization symmetry $\mathrm{Diff}(S^{1})/SL(2,\mathbb{R})$ familiar from SYK-like models. These modes exhibit an Ohmic spectrum,
$A_{\varepsilon}(\omega)\sim\frac{|\omega|}{T_{0}+T},$ (33)
which enforces analytic low-frequency behavior and furnishes an efficient energy-relaxing bath without relaxing momentum. The scale $T_{0}$ regularizes the IR and interpolates to Fermi-liquid-like behavior at $T\ll T_{0}$. To analyze transport in a system where total momentum is only weakly relaxed, it is necessary to project out the component of the electric current parallel to momentum. We therefore introduce the incoherent current operator
$J^{\mathrm{inc}}\equiv J-\frac{\chi_{JP}}{\chi_{PP}}\,P,\qquad\chi_{J^{\mathrm{inc}}P}=0,$ (34)
where $\chi_{\alpha\beta}$ are static susceptibilities. The slow operator basis is then chosen as $A_{\alpha}=\{J^{\mathrm{inc}},P\}$.
The dc conductivity follows from the standard Mori–Zwanzig memory–matrix expression
$\sigma_{ij}=\sum_{\alpha\beta}\chi_{J_{i}A_{\alpha}}\,\bigl[M(0)+N\bigr]^{-1}_{\alpha\beta}\,\chi_{A_{\beta}J_{j}},$ (35)
where $M$ is the symmetric memory matrix and $N$ encodes the antisymmetric Lorentz force, $N_{P_{x}P_{y}}=-N_{P_{y}P_{x}}=\omega_{c}\chi_{PP}$. For later convenience we introduce the shorthand
$Q\equiv J^{\mathrm{inc}},\qquad\chi_{QQ}\equiv\chi_{J^{\mathrm{inc}}J^{\mathrm{inc}}}.$ (36)
No new operator is introduced; $Q$ is simply the incoherent current. Projecting (35) onto the $\{Q,P\}$ block yields
$\sigma_{xx}^{\mathrm{inc}}=\frac{\chi_{QQ}}{\Gamma_{Q}},\qquad\rho_{xx}^{\mathrm{inc}}=\frac{\Gamma_{Q}}{\chi_{QQ}},$ (37)
with the incoherent relaxation rate
@@
The operator $\dot{Q}$ couples directly to the emergent reparametrization soft modes. Using (33) in (38), and integrating over frequencies with a thermal cutoff of order $T$, one finds
$M_{QQ}(0)\sim\int_{0}^{T}d\omega\,\frac{\omega}{T_{0}+T}\sim\frac{T^{2}}{T_{0}+T},\qquad\Rightarrow\qquad\Gamma_{Q}(T)\sim\frac{T^{2}}{T_{0}+T}.$ (39)
This relaxation rate crosses over from Fermi–liquid–like behavior $\Gamma_{Q}\sim T^{2}/T_{0}$ for $T\ll T_{0}$ to Planckian behavior $\Gamma_{Q}\sim T$ for $T\gg T_{0}$. The momentum channel is governed by lattice umklapp processes. The corresponding memory element is
@@
where $\mathbf{G}$ are reciprocal–lattice vectors and $\chi^{R}_{\rho_{\mathbf{G}}\rho_{-\mathbf{G}}}$ is the density–response function at umklapp momentum. Standard kinematic arguments lead to
$M_{PP}(0)\propto T^{2},\qquad\Gamma_{P}\equiv\frac{M_{PP}(0)}{\chi_{PP}}\propto T^{2},$ (41)
@@
independent of the soft mode physics. The separation of channels is therefore not an assumption but a consequence of the distinct operators $Q$ and $P$ entering the memory matrix. The separation of channels is therefore not an assumption but a consequence of the different operators $Q$ and $P$ entering the memory matrix. Deviations from $\cot\theta_{H}\propto T^{2}$ in electron-doped cuprates [51] and ruthenates [52] reflect orbital selectivity and multi-band scattering, but the separation of channels remains a robust organizing principle [53].